OpenStudy (anonymous):

simplify (sqrt (x))/(1+sqrt(x))

hartnn (hartnn):

to rationalize the denominator,multiply numerator and denominator by (1-sqrt x)

OpenStudy (anonymous):

\[\sqrt{x}\] over \[1+\sqrt{x}\] ?

OpenStudy (unklerhaukus):

\[\frac{\sqrt x}{1+\sqrt x}\]\[=\frac{\sqrt x}{1+\sqrt x}\times1\]\[=\frac{\sqrt x}{1+\sqrt x}\times\frac{1-\sqrt x}{1-\sqrt x}\]\[=\]

OpenStudy (anonymous):